Monica has $1, $5, and $10 bills in her wallet that are worth $96. If she
had one more $1 bill, she would have just as many $1 bills as $5 and $10
bills combined. She has 23 bills total. How many of each denomination
does she have?Please show me the steps.

Answer by xinxin(76)

Let x = # of $1 bill, y = # of $5 bill, z = # of $10 bill, then there're 3 functions you can get:
(1). x + 5y + 10z = 96
(2). x +1 = y + z
(3). x + y + z =23
From (1), (2), get 6y + 11z = 97 (4)
From (2), (3), get 2y + 2z = 24 (5)
From (4), (5), get y = 7, z =5, plug these into (2) get x = 11